# Neural Value Iteration

### v2

@@ 5.2 Offline POMDP Methods
MCVI $17.34\pm 2.52$ $13.78\pm 2.13$ $10.78\pm 1.92$ $2.36\pm 2.23$ $2.94\pm 0.15$ $0.68\pm 0.16$
NVI $18.95\pm 1.49$ $17.51\pm 1.56$ $\mathbf{13.60\pm 2.27}$ $\mathbf{11.63\pm 1.09}$ $3.73\pm 0.10$ $\mathbf{2.03\pm 0.03}$
@@ 17 $\mathcal{B}\leftarrow\mathcal{B}\cup\{b^{a,o}\}$ ;
For the planning experiments, we used the official POMCGS package POMCGraphSearch at https://github.com/ori-goals/POMCGraphSearch.jl, and implemented the baseline method MCVI in Julia using the POMDP.jl framework. We applied several custom optimizations, including vectorized operations, multi-threading, value caches, and static value types to accelerate computation. These improvements enabled faster performance compared to existing Julia implementations (e.g., https://github.com/JuliaPOMDP/MCVI.jl) in our tested domains.
@@
Network Architecture
In NVI, the POMDP value function is approximated by a finite set of neural networks, each denoted as $\alpha_{\text{NN}}$. In our experiments, we trained all $\alpha_{\text{NN}}$ with multilayer perceptrons (MLPs) with ReLU activations. The network architectures are as follows:
@@
  • Light Dark and Lidar Roomba:

### v3

@@ 5.2 Offline POMDP Methods
MCVI $17.34\pm 2.52$ $13.78\pm 2.13$ $10.78\pm 1.92$ $4.57\pm 2.19$ $2.94\pm 0.15$ $0.68\pm 0.16$
NVI $18.95\pm 1.49$ $17.51\pm 1.56$ $\mathbf{13.60\pm 2.27}$ $\mathbf{12.31\pm 1.65}$ $3.73\pm 0.10$ $\mathbf{2.03\pm 0.03}$
@@ 17 $\mathcal{B}\leftarrow\mathcal{B}\cup\{b^{a,o}\}$ ;
There are three planning algorithms compared in our experiments: SARSOP, POMCGS, and MCVI. For POMCGS, we used the official POMCGS Julia implementation available at https://github.com/ori-goals/POMCGraphSearch.jl. For MCVI, we implemented a customized version in Julia using the POMDP.jl framework. We applied several custom optimizations, including vectorized operations, multi-threading, and static value types to accelerate computation. These improvements enabled faster performance compared to existing Julia implementation (e.g., https://github.com/JuliaPOMDP/MCVI.jl) in our tested domains.
@@
Network Architecture Selection
In NVI, the POMDP value function is approximated by a finite set of neural networks, each denoted as $\alpha_{\text{NN}}$. In our experiments, we trained all $\alpha_{\text{NN}}$ with multilayer perceptrons (MLPs) with ReLU activations.
We evaluated several candidate architectures with varying depth and width to balance approximation capacity and computational efficiency. The tested configurations included:
Dense(input_dim, 64, relu)
Dense(64, 64, relu)
Dense(64, 1)
Dense(input_dim, 128, relu)
Dense(128, 64, relu)
Dense(64, 32, relu)
Dense(32, 1)
Dense(input_dim, 128, relu)
Dense(128, 64, relu)
Dense(64, 64, relu)
Dense(64, 1)
Dense(input_dim, 256, relu)
Dense(256, 128, relu)
Dense(128, 64, relu)
Dense(64, 1)
Dense(input_dim, 512, relu)
Dense(512, 256, relu)
Dense(256, 256, relu)
Dense(128, 64, relu)
Dense(64, 1)
Empirically, increasing network size beyond a moderate capacity did not produce noticeable improvements in policy quality. In contrast, smaller MLPs achieved comparable performance while reducing computational cost. In practice, the training time per backup was typically below 5 seconds, including CPU–GPU data transfer overhead with small MLPs.
Based on this empirical trade-off, we selected the following architectures:
@@
  • Light-Dark and Lidar Roomba:
@@
Input and Output of Network
For each $\alpha_{\text{NN}}$, the input is a vectorized state representation and the output is a scalar, consistent with the definition of an $\alpha$-vector as a mapping from $\mathbb{S}$ to $\mathbb{R}$. Specifically, in RockSample, the input consists of a one-hot encoding of the robot position concatenated with the rock status vector. In LightDark and LidarRoomba, the input is the state vector representing the agent’s location.